OpenStudy (anonymous):

The function H described by H(x) =2.75x + 71.48 cab be used to predict the height in cm, of a womans whose humerus is x cm long. Predict the height of a woman whose humerus is 32 cm long

OpenStudy (anonymous):

plug x= 32 and solve

OpenStudy (anonymous):

nee help?

OpenStudy (anonymous):

H(32) =2.75(32) + 71.48 = 88+71.48 =159.49

OpenStudy (anonymous):

If I deserve medal then give me one